The boiling point of 200 g of H2O containing 30.0 g of a non-ionic substance is 101.4 degrees C. What is the molecular weight of the non-ionic substance?

Kf(water) = 1.86

Kb(water)= 0.51
Answer #1

The molality of the solution is calculated, with the equation of boiling point increase:

m = ΔT / K = (101.4 - 100) / 0.51 = 2.75 m

The moles of the solute are calculated:

n = m * Kg solvent = 2.75 m * 0.2 Kg = 5.5 mol

The molar weight of the solute is calculated:

MM = g / n = 30 g / 5.5 mol = 5.45 g / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.